# Adaptive Coverage Policies in Conformal Prediction

### v1

@@ 1 Introduction
A central limitation of classical conformal prediction is the need to specify a fixed coverage level $1-\alpha$. This choice is typically arbitrary and has received relatively little attention in the literature, despite its significant impact on the size and informativeness of the prediction sets. In practice, small values of $\alpha$ can yield overly conservative prediction sets, while large values may lead to empty sets that are uninformative. Moreover, classical conformal methods do not allow $\alpha$ to be chosen based on the observed data, because the marginal coverage guarantee (1) only holds when $\alpha$ is fixed before seeing the calibration or test points. In some problems, a data analyst might prefer to adjust $\alpha$ after inspecting the data. For example, suppose that $\alpha=1\%$ yields an extremely large prediction set, while $\alpha=1.1\%$ gives a much more compact set with only a slight relaxation in coverage. Most practitioners would naturally opt for the latter, as it provides a more informative and usable prediction. Unfortunately, in traditional conformal prediction, choosing $\alpha$ after inspecting the data undermines the marginal coverage guarantee (1) and is closely related to the phenomenon of p-hacking in the statistical literature (Simmons et al., 2011; Head et al., 2015). This rigidity is limiting in settings where uncertainty varies across examples, or where we may want to tailor coverage to the difficulty of each instance.
@@
We tackle this challenge by leveraging recent advances in e-values (Shafer and Vovk, 2019; Vovk and Wang, 2021; Grünwald et al., 2024; Ramdas and Wang, 2024) and post-hoc inference (Wang and Ramdas, 2022; Xu et al., 2024; Grünwald, 2024; Koning, 2024; Gauthier et al., 2025b; Chugg et al., 2025), which provide valid coverage guarantees even when the miscoverage level $\alpha$ is selected adaptively. Building on this, we propose to optimize the choice of $\alpha$ by training a neural network on observed data. To construct training examples, we adopt a leave-one-out approach on the calibration set: for each held-out point, the remaining points serve as a pseudo calibration set, and the held-out point acts as a pseudo test sample. This generates a collection of labeled examples, enabling a model to predict a sample-specific coverage policy. Our aim is to design a mapping from each pseudo calibration-test pair to a coverage policy that maximizes the informativeness of the resulting prediction sets. Crucially, this approach allows the data-dependent miscoverage level $\tilde{\alpha}$ to adapt to individual test samples while maintaining valid marginal guarantees, offering a more flexible and data-driven alternative to choosing a fixed $\alpha$ in conventional conformal prediction.
By leveraging the post-hoc validity of e-values, this approach enables adaptive conformal prediction, modulating prediction set size according to instance difficulty. Our work is related to recent efforts that combine e-values with conformal prediction (Vovk, 2025; Gauthier et al., 2025b). The leave-one-out strategy we use has also been employed in prior work (Gauthier et al., 2025a), though in a different role—there it was introduced to estimate marginal adaptive miscoverage, whereas here we employ the strategy to generate training data for fitting an adaptive coverage policy.
@@ 2.1 Conformal e-prediction
While conformal e-prediction is compatible with any valid e-variable, for concreteness we employ the so-called soft-rank e-variable. This construction, which first appeared in Wang and Ramdas (2022) and Koning (2025), and was later applied in the context of conformal prediction by Balinsky and Balinsky (2024), takes the following form:
@@ 2.2 Post-hoc validity
Consider a calibration set $\{(X_{i},Y_{i})\}_{i=1}^{n}$ and a test data point $(X_{\rm test},Y_{\rm test})$ such that $(X_{1},Y_{1}),\dotsc,(X_{n},Y_{n}),(X_{\rm test},Y_{\rm test})$ are exchangeable. Let $\tilde{\alpha}>0$ be any miscoverage level that may depend on this data. Then we have that:
@@
Proposition 2.2 enables us to obtain marginal guarantees for any coverage level, including those that depend on the data. Building on this flexibility, we aim to design a coverage policy that adapts the miscoverage level in order to minimize the size of the resulting prediction sets. Informally, a coverage policy is simply a rule that maps the calibration scores and potential test scores to a (data-dependent) miscoverage level $\tilde{\alpha}\in(0,1)$. We will provide a formal definition in Definition 2.5, but before doing so, we first examine how prediction set sizes behave in both the classification and regression settings. This detour is useful since we want to design coverage policies that minimize these sizes, and the form they take in these two cases will guide us toward a simplified definition of coverage policies.
@@
Given a calibration set $(X_{i},Y_{i})_{i=1}^{n}$ and a test feature $X_{\rm test}$, in classification problems the size of a conformal set at miscoverage level $\tilde{\alpha}\in(0,1)$ is given by
$\rm Size\big(\hat{C}_{n}^{\tilde{\alpha}}(X_{\text{test}})\big)=\#\left\{y\in\mathcal{Y}:\frac{S(X_{\rm test},y)}{\frac{1}{n+1}\left(\sum_{i=1}^{n}S(X_{i},Y_{i})+S(X_{\rm test},y)\right)}<\frac{1}{\tilde{\alpha}}\right\}$.
@@
$\sum_{y\in\mathcal{Y}}\mathbb{1}{\left\{\frac{S(X_{\rm test},y)}{\tfrac{1}{n+1}\left(\sum_{i=1}^{n}S(X_{i},Y_{i})+S(X_{\rm test},y)\right)}\!<\!\frac{1}{\tilde{\alpha}}\right\}}\!\approx\!\sum_{y\in\mathcal{Y}}\sigma\!\left(k\!\left(\frac{1}{\tilde{\alpha}}-\frac{S(X_{\rm test},y)}{\tfrac{1}{n+1}\left(\sum_{i=1}^{n}S(X_{i},Y_{i})+S(X_{\rm test},y)\right)}\right)\!\right),$
@@
that maps the sum of calibration scores and a test summary statistic to a miscoverage level. [^1] While we refer to it as a coverage policy, the function in fact outputs a miscoverage level. This is convenient, since the conformal set is obtained directly from the inverse of the miscoverage. Here, $\mathcal{T}$ denotes the space of possible test statistics $t(X_{\rm test})$: in classification, $t(X_{\rm test})$ is the vector of potential scores $(S(X_{\rm test},y))_{y\in\mathcal{Y}}$, whereas in regression it is empty, reflecting that the conformal set size (5) does not depend on the test feature. The output
@@ 2.4 Selecting $\lambda$: insights from the constant-$\alpha$ setting
The choice of the regularization term $\lambda$ is particularly critical in practice, as it directly impacts, among other things, the expected size of the prediction set at test time. One practical strategy to select $\lambda$ is to monitor the behavior of the network $\tilde{\alpha}_{\theta}$ during training and track the final values of the average sizes of the training conformal sets.
@@
Theorem 2.6, proved in Appendix A, shows that in the idealized case where the network output is constant, the average size of the conformal sets provides a reliable estimate of the expected size at test time. This result serves as a theoretical foundation for our approach: by monitoring the leave-one-out sizes during training, we can approximate the expected test-time behavior of the conformal predictor.
@@
A key ingredient in this procedure is the monotonicity of the leave-one-out size with respect to $\lambda$. In the constant-$\alpha$ setting, Proposition 2.7, proved in Appendix B, establishes this property formally. Monotonicity ensures that we can safely use a bracketing-and-bisection strategy: once we find two values of $\lambda$ such that the leave-one-out average lies below and above $M$, repeated halving of the interval will converge to the desired value. Empirically, we observe that this monotonicity extends to the trained, input-dependent setting, which makes the bracketing-and-bisection procedure effective in practice.
@@ 3 Experimental evaluation
For a calibration set sampled uniformly at random, Figure 1 illustrates the training dynamics of the coverage policy produced by Algorithm 1 under different choices of the regularization strength $\lambda~\in~\{5,10,50\}$. The training curves decrease and converge, indicating that the model is optimizing effectively. Moreover, both the prediction set size and the expected miscoverage stabilize asymptotically as training progresses. This effective convergence highlights the value of the leave-one-out procedure, which provides sufficient signal for successfully training a coverage policy.
@@ Acknowledgements
The first term is a standard concentration term. Since the $S(X_{i},Y_{i})$ are i.i.d. and bounded by assumption, Hoeffding’s inequality [Hoeffding, 1963] gives, for any $\delta>0$,
@@
We now turn to case (i). The argument follows the strategy of Gauthier et al. [2025a, Theorem 3.1]. We use the same notations: for each calibration index $j=1,\dots,n$, define the random vector
@@
for all $\delta>0$, where the last inequality follows from Gauthier et al. [2025a]. By Hoeffding’s inequality,
@@
again using Gauthier et al. [2025a]. Combining the three terms with an union bound, we conclude that with probability $\geq 1-\delta$,
@@ Appendix D Complexity analysis and resource details
In the classification setting with the loss (2.3), the time complexity of Algorithm 1 is

### v2

@@ 1 INTRODUCTION
A core feature of classical conformal prediction is the specification of a fixed coverage level $1-\alpha$ prior to observing the data. While this provides a highly desirable and easily interpretable statistical guarantee, it imposes a rigid operational constraint. In practice, this fixed choice can lead to uninformative predictions: small values of $\alpha$ can yield overly conservative prediction sets, while large values may lead to empty sets. Moreover, classical conformal methods do not allow $\alpha$ to be chosen based on the observed data, because the marginal coverage guarantee (1) only holds when $\alpha$ is fixed before seeing the calibration or test points. In many operational settings, a practitioner might prefer to adjust $\alpha$ after inspecting the data. For instance, in a medical diagnosis task, a standard conformal method guaranteeing 99% coverage might output a broad set of seven diagnoses. While statistically valid, this set may be clinically unactionable. A physician might prefer to dynamically relax the coverage slightly (e.g., from $\alpha=1\%$ to $\alpha=2\%$) to yield a precise, actionable set of three diagnoses. Most practitioners would naturally opt for the latter, as it provides a more informative and usable prediction. Unfortunately, in traditional conformal prediction, choosing $\alpha$ after inspecting the data undermines the marginal coverage guarantee (1) and is closely related to the phenomenon of p-hacking in the statistical literature (Simmons et al., 2011; Head et al., 2015). This rigidity is limiting in settings where uncertainty varies across examples, or where we may want to tailor coverage to the difficulty of each instance.
@@
We tackle this challenge by leveraging recent advances in e-values (Shafer and Vovk, 2019; Vovk and Wang, 2021; Grünwald et al., 2024; Ramdas and Wang, 2025) and post-hoc inference (Wang and Ramdas, 2022; Xu et al., 2024; Grünwald, 2024; Koning, 2024; Gauthier et al., 2025b; Chugg et al., 2026), which provide valid coverage guarantees even when the miscoverage level $\alpha$ is selected adaptively. It is known that conformal sets constructed using e-values generally yield slightly larger prediction sets than those built with standard p-values for a given fixed $\alpha$ (Vovk, 2025). However, e-values are necessary to enable this adaptivity as they are the only way to construct post-hoc p-values, the statistical objects that allow for valid inference even when the miscoverage level is chosen after observing the data (Koning, 2024). We willingly trade a slight loss in fixed-$\alpha$ efficiency for this powerful operational advantage, as it allows the coverage level to be selected adaptively without invalidating the statistical guarantees.
Building on this, we propose to optimize an adaptive coverage policy by training a neural network on observed data. To construct training examples, we adopt a leave-one-out approach on the calibration set: for each held-out point, the remaining points serve as a pseudo calibration set, and the held-out point acts as a pseudo test sample. This generates a collection of labeled examples, enabling a model to predict a sample-specific coverage policy. Our aim is to design a mapping from each pseudo calibration-test pair to a coverage policy that maximizes the informativeness of the resulting prediction sets. We formalize this informativeness as a dual objective (detailed in Section 2.3): minimizing the prediction set size while simultaneously minimizing the adaptive miscoverage level $\tilde{\alpha}$. This trade-off is governed by a user-specified regularization parameter $\lambda$. Furthermore, in Section 2.4, we introduce a principled procedure to select $\lambda$ such that the expected prediction set size at test time meets a desired target. Crucially, this approach allows the data-dependent miscoverage level $\tilde{\alpha}$ to adapt to individual test samples while maintaining valid marginal guarantees, offering a more flexible and data-driven alternative to choosing a fixed $\alpha$ in conventional conformal prediction.
By leveraging the post-hoc validity of e-values, this approach enables adaptive conformal prediction, modulating prediction set size according to instance difficulty. Our work is related to recent efforts that combine e-values with conformal prediction (Vovk, 2025; Gauthier et al., 2025b). In particular, we contrast our approach with recent prior work (Gauthier et al., 2025a; Liu et al., 2026), which also utilizes e-values and a leave-one-out strategy. However, our objective differs fundamentally: while that method inverts the conformal procedure to find the minimum $\alpha$ satisfying a hard size constraint for a specific prediction set, our method learns a parametric coverage policy to optimize a soft constraint by targeting an expected prediction set size at test time. Furthermore, the two methods employ the leave-one-out procedure for entirely different purposes. Whereas the prior work uses it solely to estimate coverage guarantees, we utilize it as an integral generative step to construct the training data required to fit our neural coverage policy.
@@ 2.1 Conformal e-prediction
While conformal e-prediction is compatible with any valid e-variable, for concreteness we employ the soft-rank e-variable. This construction, which first appeared in Wang and Ramdas (2022) and Koning (2025), and was later applied in the context of conformal prediction by Balinsky and Balinsky (2024), takes the following form:
@@ 2.2 Post-hoc Validity
Consider a calibration set $\{(X_{i},Y_{i})\}_{i=1}^{n}$ and a test data point $(X_{\rm test},Y_{\rm test})$ such that $(X_{1},Y_{1}),\dotsc,(X_{n},Y_{n}),(X_{\rm test},Y_{\rm test})$ are exchangeable. Let $\tilde{\alpha}>0$ be any miscoverage level may depend on all of these data points. Then we have that:
@@
Proposition 2.2 enables us to obtain marginal guarantees for any coverage level, including those that depend on the data. This post-hoc validity holds because the definition of the e-value relies strictly on the non-conformity scores, which are exchangeable by assumption. Because an adaptive policy does not alter the computation of these underlying scores, selecting the miscoverage level $\tilde{\alpha}$ in a data-dependent manner does not retroactively compromise exchangeability. Building on this flexibility, we aim to design a coverage policy that adapts the miscoverage level in order to minimize the size of the resulting prediction sets. Informally, a coverage policy is simply a rule that maps the calibration scores and potential test scores to a (data-dependent) miscoverage level $\tilde{\alpha}\in(0,1)$. We will provide a formal definition in Definition 2.5, but before doing so, we first examine how prediction set sizes behave in both the classification and regression settings. This detour is useful since we want to design coverage policies that minimize these sizes, and the form they take in these two cases will guide us toward a simplified definition of coverage policies.
@@
Given a calibration set $(X_{i},Y_{i})_{i=1}^{n}$ and a test feature $X_{\rm test}$, in classification problems the size of a conformal set $\rm Size\big(\hat{C}_{n}^{\tilde{\alpha}}(X_{\text{test}})\big)$ at miscoverage level $\tilde{\alpha}\in(0,1)$ is given by
$\left\{y\in\mathcal{Y}:\frac{S(X_{\rm test},y)}{\tfrac{1}{n+1}\!\left(\sum_{i=1}^{n}S(X_{i},Y_{i})+S(X_{\rm test},y)\right)}\!<\!\frac{1}{\tilde{\alpha}}\right\}$.
@@
$\sum_{y\in\mathcal{Y}}\mathbb{1}{\left\{\frac{S(X_{\rm test},y)}{\frac{1}{n+1}\left(\sum_{i=1}^{n}S(X_{i},Y_{i})+S(X_{\rm test},y)\right)}<\frac{1}{\tilde{\alpha}}\right\}}$
$\!\approx\!\sum_{y\in\mathcal{Y}}\!\sigma\!\left(\!k\!\left(\frac{1}{\tilde{\alpha}}\!-\!\frac{S(X_{\rm test},y)}{\tfrac{1}{n+1}\left(\sum_{i=1}^{n}S(X_{i},Y_{i})+S(X_{\rm test},y)\right)}\!\right)\!\right),$ (5)
@@
that maps the sum of calibration scores and a test summary statistic to a miscoverage level. [^1] While we refer to it as a coverage policy, the function actually outputs a miscoverage level $\tilde{\alpha}$. This convention simplifies our notation, since the conformal set is defined directly using the inverse $1/\tilde{\alpha}$. Here, $\mathcal{T}$ denotes the space of possible test statistics $t(X_{\rm test})$: in classification, $t(X_{\rm test})$ is the vector of potential scores $(S(X_{\rm test},y))_{y\in\mathcal{Y}}$, whereas in regression it is empty, reflecting that the conformal set size (6) does not depend on the test feature. The output
@@ 2.4 Selecting $\lambda$: Insights From the Constant-$\alpha$ Setting
The choice of the regularization term $\lambda$ is critical in practice, as it directly impacts, among other things, the expected size of the prediction set at test time. One practical strategy to select $\lambda$ is to monitor the behavior of the network $\tilde{\alpha}_{\theta}$ during training and track the final average sizes of the training conformal sets.
@@
Theorem 2.6, proved in Appendix A, shows that in the idealized constant-output case, the average size of the conformal sets provides a reliable estimate of the expected size at test time. This result provides a theoretical foundation for our approach: monitoring the leave-one-out sizes during training to approximate the expected test-time behavior of the conformal predictor.
@@
A key ingredient in this procedure is the monotonicity of the leave-one-out size with respect to $\lambda$. In the constant-$\alpha$ setting, Proposition 2.7, proved in Appendix B, establishes this property formally. Monotonicity ensures that we can safely use a bracketing-and-bisection strategy: once we identify two values of $\lambda$ such that the leave-one-out average lies below and above $M$, repeated halving of the interval guarantees convergence to the desired value. Empirically, we observe that this monotonicity approximately holds in the trained, input-dependent setting, ensuring the reliability of the bracketing-and-bisection procedure in practice.
@@ 3 EXPERIMENTAL EVALUATION
For a calibration set sampled uniformly at random, Figure 1 illustrates the training dynamics of the coverage policy produced by Algorithm 1 under different choices of the regularization strength $\lambda\in\{5,10,50\}$. The training curves decrease and converge, indicating that the model is optimizing effectively. Moreover, both the prediction set size and the expected miscoverage stabilize asymptotically as training progresses. This effective convergence highlights the value of the leave-one-out procedure, which provides sufficient signal for successfully training a coverage policy.
@@ Acknowledgements
The authors thank the anonymous reviewers for their helpful feedback that improved this work.
@@
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [No] We will publicly release the code.
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes]
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes]
    2. (b)
      The license information of the assets, if applicable. [Yes]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Yes]
    4. (d)
      Information about consent from data providers/curators. [Yes]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
Supplementary Materials
@@
The first term is a standard concentration term. Since the $S(X_{i},Y_{i})$ are i.i.d. and bounded by assumption, Hoeffding’s inequality (Hoeffding, 1963) gives, for any $\delta>0$,
@@
We now turn to case (i). The argument follows the strategy of Gauthier et al. (2025a, Theorem 3.1). We use the same notations: for each calibration index $j=1,\dots,n$, define the random vector
@@
for all $\delta>0$, where the last inequality follows from Gauthier et al. (2025a). By Hoeffding’s inequality,
@@
again using Gauthier et al. (2025a). Combining the three terms with an union bound, we conclude that with probability $\geq 1-\delta$,
@@ Appendix C ADDITIONAL EXPERIMENTS
Finally, we note an important practical limitation of this specific regression setup. As derived in Remark 6, the conformal set size for the MAE score using the soft-rank e-variable does not depend on the test feature $X_{\rm test}$. Consequently, the resulting prediction sets maintain a constant width across the feature space and cannot adapt to varying levels of uncertainty or noise in the data. While this makes the sets overly large and rigid for practical, real-world regression tasks, this setup serves as a clean, differentiable testbed to empirically validate our procedure. For practical regression applications, combining our adaptive coverage framework with locally adaptive non-conformity scores or alternative e-variable constructions would be more appropriate.
@@ Appendix D COMPLEXITY ANALYSIS AND RESOURCE DETAILS
In our complexity analysis, we assume that the base model $f$ is a frozen black box. Therefore, the non-conformity scores $S(X_{i},y)$ for all calibration points and all classes can be pre-computed prior to training. Once these scores are cached, querying $S$ inside the training loop becomes a constant-time memory lookup. Under this pre-computation, in the classification setting with the loss (2.3), the time complexity of the optimization loop in Algorithm 1 is
@@
To address the scalability of our approach to domains substantially larger than CIFAR-10 (such as ImageNet), we note that Algorithm 1 operates exclusively on the pre-computed non-conformity scores. It is entirely independent of the raw input dimensionality (e.g., image resolution). As established above, the training complexity scales linearly with the calibration set size $n$ and the number of classes $|\mathcal{Y}|$. Even for massive datasets like ImageNet, the required calibration set size $n$ typically remains small and manageable (e.g., $n\approx 1000$). Thus, for a large-scale task ($n=1000$, $|\mathcal{Y}|=1000$), the neural coverage policy $\tilde{\alpha}_{\theta}$ is simply trained over low-dimensional score vectors (of size 1001). The linear scaling in $n$ and $|\mathcal{Y}|$: adds negligible computational overhead. The total wall-clock time is therefore overwhelmingly dominated by the initial forward passes of the base predictor $f$ needed to cache the scores, ensuring our adaptive conformal wrapper remains highly scalable to large real-world tasks on standard hardware.